A child does 500 J of work while pulling a box from the ground up to his tree house at a steady rate of 2 m/s2 with a light rope. The tree house is 4 m above the ground. What is the mass of the box?
Answer #1

Solution)

We know, work done=mgh

Where, work done, W=500 J

Acceleratio, a=2m/s^2

Height, h=4 m

So,

500 J=m*2*4

Mass, m=500/8= 62.5 kg

========